OpenStudy (anonymous):

What is the solution to the equation? 5x + 10 = –4x – 17 A. B. C. D.

OpenStudy (anonymous):

sorry options wont load can you just solve

OpenStudy (amoodarya):

put variables in left side and numbers in right side simplify then you have have the answer

OpenStudy (amoodarya):

5x + 10 = –4x – 17 5x+4x=-17-10 9x=-27 x=-3